Please help. I'm not sure what to debit and what to credit?Question 9 0.9 pts Presented below are selected account balances for Pitchfork, Inc. as of December 31, X1. Using this inform
Answer #1

Answer:

The required closing entry to move net income into its permanent equity account is as follows:

Debit Income Summary , for $73,300 Credit Retained Earnings arnings 1. for I $73.300

Working:

Net Sales Revenue = Sales Revenue - Sales Return - Sales discount = 410000 - 12000 - 15000 = $383,000

Total expenses = Cost of Goods sold + Selling Expenses + Administrative expense + Income tax expense

=225700 + 16000 + 38000 + 30000

= $309,700

Net Income = Net sales revenue - Total expenses = 383000 - 309700 = $73,300

The T account of Income Summary account on posting the closing entry to close revenue to Income Summary and the closing entry to close expenses to Income Summary will be as follows:

Income Summary Account Dec 31, X1 $309,700 $383,000 Dec 31, X1 $73,300 Balance

As such after posting the two closing entries as above, the Income summary account will have a credit balance of $73,300, which is the net income.

Next closing entry to close Income Summary account and transfer net income to retained earnings will be:

Debit Income Summary $73,300

Credit Retained Earnings $73,300